OpenStudy (anonymous):

Write the equation 5x - y + 6 = 0 in standard form with integer coefficients. @Orion1213 can u help meh with this pls

OpenStudy (anonymous):

is it Y=5X+6

OpenStudy (superhelp101):

uh i got 5x-y=-6

OpenStudy (superhelp101):

ohhhhh nvm

OpenStudy (superhelp101):

u are right! :) good job! :)

OpenStudy (anonymous):

thx can u help meh with more ques i have a test

OpenStudy (superhelp101):

yeah! just post it on a different thread plz! :D

OpenStudy (anonymous):

ok just a sec

OpenStudy (texaschic101):

standard form is Ax + By = C so it would be : 5x - y = -6 or 5x - 1y = -6